# Colorful Minors

### v1

@@ Abstract
We introduce the notion of colorful minors, which generalizes the classical concept of rooted minors in graphs. A -colorful graph is defined as a pair where is a graph and assigns to each vertex a (possibly empty) subset of at most colors. The colorful minor relation enhances the classical minor relation by merging color sets at contracted edges and allowing the removal of colors from vertices. This framework naturally models algorithmic problems involving graphs with (possibly overlapping) annotated vertex sets. We develop a structural theory for colorful minors by establishing several theorems characterizing -colorful minor-free graphs, where consists either of a clique or a grid with all vertices assigned all colors, or of grids with colors segregated and ordered on the outer face. Leveraging our structural insights, we provide a complete classification – parameterized by the number of colors – of all colorful graphs that exhibit the Erdős–Pósa property with respect to colorful minors.
On the algorithmic side, we provide a fixed-parameter tractable algorithm for colorful minor testing and a variant of the -disjoint paths problem. Together with the fact that the colorful minor relation forms a well-quasi-order, this implies that every colorful minor-monotone parameter on colorful graphs admits a fixed-parameter algorithm. Furthermore, we derive two algorithmic meta-theorems (AMTs) whose structural conditions are linked to extensions of treewidth and Hadwiger number on colorful graphs. Our results suggest how known AMTs can be extended to incorporate not only the structure of the input graph but also the way the colored vertices are distributed in it.
@@
- 1 Introduction
- 2 Preliminaries
- 3 Excluding a rainbow clique
- 4 Excluding a rainbow-grid
- 5 Excluding segregated grids
- 6 The Erdős-Pósa property for -colorful graphs
- 7 A colorful minor algorithm
- 8 Algorithmic Applications
- 9 Further research directions
- A Some problems where our algorithmic results apply
- B Proof of lemma 7.6
@@ 1 Introduction
Over the last decades, a prominent interplay between structural graph theory and theoretical computer science has been the search for structural properties of graphs that facilitate the design of efficient algorithms for otherwise NP-hard problems (see for example [17]). One of the most studied such parameters is a measure for the tree-likeness known in its most popular form as treewidth [75, 4, 5, 6]. Treewidth, denoted by plays an exceptional role not only in the world of (parameterized) algorithms, but also in the theory of graph minors by Robertson and Seymour [73, 76, 77, 78]. A central theorem of Robertson and Seymour is the Grid Theorem [76] that states that any minor-closed graph class has bounded treewidth if and only if does not contain all planar graphs. This reveals a fundamental limitation in the use of treewidth (or any minor-monotone parameter really) as a parametrization for an NP-hard problem: While there exists a wide range of problems that are tractable on classes of bounded treewidth, there exist many important problems that are NP-hard even on planar graphs [33, 16]. This means that, in the regime of minor-closed graph classes, such problems are tractable if and only if a planar graph is excluded, unless PNP.
@@
A particularly interesting class of problems tractable on graphs of bounded treewidth but NP-hard on a more general proper minor-closed graph class are certain problems whose instances come with specific vertex sets, typically called roots, terminals, or annotated vertices. A prototypical example of such a problem is Steiner Tree. The Steiner Tree problem receives a graph accompanied with a special set of annotated vertices as input and asks for the minimum number of edges in a connected subgraph of containing all vertices of The problem is known to be NP-complete on planar graphs, but it can be solved in polynomial time when the relative structure of the set is further restricted [20, 23, 54]. There exists a variety of problems exhibiting similar properties [32, 16, 60, 67].
To formalize the above notion of relative structure one often uses the concept of so-called rooted minors. A pair where is a graph and is called an annotated graph. An annotated graph is a rooted minor of an annotated graph if there exists a collection of pairwise vertex-disjoint connected subgraphs of such that for all there is an edge between and and for all we have
@@
Computational problems such as Steiner Tree can be naturally understood as problems on annotated graphs. Recently, several positive algorithmic results for Steiner Tree on annotated graphs excluding certain types of rooted minors have emerged [41, 45] alongside purely structural results on rooted minors [64, 44, 28] and found various applications [44, 2]. The study of Erdős-Pósa dualities lies at the crossroads of structural and algorithmic graph theory. The Erdős-Pósa property of (rooted) minors asks for a fixed (annotated) graph if for any (annotated) graph there are either many vertex-disjoint occurrences of as a (rooted) minor, or has a small vertex set whose deletion removes all occurrences of as a (rooted) minor. Note that not every (annotated) graph has the Erdős-Pósa property. The Erdős-Pósa property has been investigated by many authors in many settings in the past [76, 68] and several key cases are known also for rooted minors [48, 70, 61, 44]. Finally, the bidimensionality or annotated treewidth of a vertex set in a graph denoted by is the largest such that contains the -grid with all vertices annotated as a rooted minor [82]. It was recently shown that a reasonable extension of Courcelle’s Theorem [15, 13, 14] exists for every minor-closed graph class when formulas are restricted to quantify over vertex-sets of small bidimensionality [81].
@@
While the concept of rooted minors in annotated graphs seems to be relatively well explored, a general theory unifying the results above appears to be lacking. While many results above are inspired by the theory of graph minors, no systematic extension of its powerful toolkit to rooted or multi-annotated minors has yet emerged. This is surprising given the great structural intrigue and algorithmic potential of these concepts as we lay out below.
@@ 1.1 Our results
Main combinatorial results.
Our core contributions are a series of structural theorems for -colorful minor-free -colorful graphs, both for general and for the particular case where is planar – under two distinct assumptions on the distribution of colors. Building on these structural results, we derive a complete classification of all -colorful graphs that satisfy the Erdős-Pósa property, for every (under the colorful minor relation).
A colorful link between algorithms and combinatorics.
Our next step is to apply our structural theorems to algorithmic settings, with the broader goal of developing an algorithmic graph structure theory applicable to problems involving vertex annotations. To this end, we first observe that colorful graphs are well quasi-ordered under the colorful minor relation. We provide an algorithm that decides whether a -colorful graph is a colorful minor of a -colorful graph in time111By we mean that there is a function such that Also, given a graph we use and for the number of its vertices and edges respectively. These two facts imply that membership for any colorful minor-closed class of colorful graphs is decidable in polynomial time.
Next, we show how classical algorithmic meta-theorems – such as Courcelle’s Theorem and the recent algorithmic meta-theorem of Sau, Stamoulis, and Thilikos [81] – can be extended to colorful graphs, where the structural conditions are expressed via the exclusion of colorful minors. This framework enables the resolution of problems on graphs containing arbitrarily large grid or clique minors.
Our results highlight a key insight: problem complexity depends not only on the graph’s structure but also on how annotations are distributed within it.
The algorithmic results presented here are not intended to be exhaustive; rather, they serve to illustrate this principle and to lay the groundwork for future exploration of algorithmic problems involving vertex annotations. In particular, we provide algorithmic applications for two of our three main structural results – Theorems 1.5 and 1.6 – while leaving the investigation of algorithmic consequences of theorem 1.7 to future work.
@@
Let be a non-negative integer. A -colorful graph is a pair where is a graph and222For we denote by and by For a vertex we call the palette of For illustrations, we use magenta for blue for , yellow for and green for
@@
A -colorful graph where is called restricted and a -colorful graph with is called empty. Note that any restricted -colorful graph is also a -colorful graph, so the notion of being restricted is highly dependent on the context given by the total set of colors. Finally, a -colorful graph is said to be rainbow333Notice that – in slight violation of the English language – we use the word “rainbow” as an adjective here. if for all
@@
See figure 1 for an example. Notice that -colorful graphs are exactly annotated graphs and the notions of colorful minors and rooted minor coincide on annotated graphs.
In section 2.4 we derive the following, as a simple consequence of a deep theorem of Robertson and Seymour [74].
@@
Fundamental algorithmic results.
One of the ground-breaking results of Robertson and Seymour’s Graph Minors Series is the Graph Minor Algorithm that allows to check if a graph is a minor of a graph in time [79] (see [53, 56] for improvements). This algorithm also solves a rooted version of the Minor Checking problem which, as a special case, contains the so-called -Disjoint Paths problem. In -Disjoint Paths the input is a graph and vertex pairs and the question is if there exist pairwise (internally) vertex-disjoint paths in such that has endpoints and for each However, the rooted version of the Graph Minor Algorithm is not sufficient to efficiently check for colorful minors even in -colorful graphs since it requires the number of roots – i.e. the number of colored vertices – to be bounded.
A fundamental application of our techniques is an adaptation of the ideas from [79] that allows for the design of a new so-called irrelevant vertex rule in the presence of a large clique minor in a -colorful graph. This allows us to efficiently solve -Colorful Minor Checking.
Theorem 1.2.
In fact, we are able to solve a -colorful rooted version of -Colorful Minor Checking that implies an efficient algorithm for a colorful generalization of -Disjoint Paths as follows.
Let be a -colorful graph, and be a sequence of (possibly empty) subsets of A -connector for and in is a tree such that
i)
both and are leaves of and, given that is the --path in
ii)
there exist distinct internal vertices of and pairwise vertex-disjoint subtrees of such that (a) appear on in the order listed, (b) for every and (c) for every
The Weak -Colorful -Disjoint Paths (-WCDP) problem has as input two integers and a -colorful graph (not necessarily distinct) vertices called the terminals, and (possibly empty) sequences called the signatures. For each there exists and such that The task is to determine whether there exist pairwise internally vertex-disjoint trees in such that for each is a -connector for and in We call the number the complexity of an instance of the Weak -Colorful -Disjoint Paths problem. Intuitively, for each one may think of the signature as the “colored sources” of systems of rivers that flow into the --path in the order listed.
Theorem 1.3.
There exist an algorithm that solves -WCDP in time given an instance of complexity on a -colorful graph
One may consider the more natural variant of the above problem asking for vertex-disjoint paths between fixed terminals, each containing specified colors in a prescribed order. Addressing this, requires the stronger framework of labeled minors, introduced by Bruhn, Joos, and Schaudt [10], which is considerably more challenging than the colorful minor setting.
Theorems 1.1 and 1.2 imply that the recognition problem for any proper colorful minor-closed class can be solved in polynomial time. We present below a parametric variant of this, based on the general framework introduced by Fellows and Langston [25, 26] for the existence of fixed-parameter tractable algorithms.
A (-colorful) graph parameter is a map assigning a non-negative integer to every (-colorful) graph. We say that a -colorful graph parameter is colorful minor monotone if for every -colorful graph and every colorful minor of we have Again, combining Theorems 1.1 and 1.2 we can deduce the following.
Theorem 1.4.
@@
We prove a total of three structure theorems as follows. The first two exclude the -colorful rainbow -clique and the -colorful rainbow -grid respectively as a colorful minor. Moreover, we generalize a theorem of Marx, Seymour, and Wollan [64], that describes the structure of -colorful graphs excluding a -grid where exactly the vertices of the first row receive color to an arbitrary number of colors.
We begin with our simplest result – excluding a rainbow clique – which can be naturally phrased in terms of a modulator to restricted graphs. For a vertex set in a graph we define the torso of in denoted by as the graph obtained from by turning the neighborhood of in into a clique for every component of In short, every -colorful graph that excludes a rainbow clique as a colorful minor has a set whose torso excludes a minor and where every component of is missing a (not necessarily the same) color.
@@
Our next theorem generalizes a theorem of Marx, Seymour, and Wollan [64] (see also [44]). Let be an integer. A -segregated grid is a -colorful graph where is the -grid where the vertices of the first column can be numbered as in order of their appearance, and we have for all and there exists a permutation of such that for every In this situation, we say that realizes . See figure 2 for an example.
Theorem 1.5 explains that every -colorful minor-free colorful graph can be turned into a collection of restricted graphs by deleting an -minor-free subgraph. Segregated grids capture exactly the situation where we require to be a planar graph. We may express this in form of a parameter as follows. The restrictive treewidth of a -colorful graph denoted by is the smallest integer such that there exists a set such that the treewidth of the torso of in is at most and for every component of the -colorful graph is restricted.
@@
contains some -segregated grid as a colorful minor and , or
@@
Notice that theorem 1.6 is optimal in the sense that on segregated grids is lower bound by a function of their size (see lemma 5.4). In other words, theorem 1.6 can be seen as a mix-max duality theorem for segregated grids. We remark that for the case has been defined by Jansen and Swennenhuis in [45] under the notation (where ).
It is easy to see that theorem 1.6 is far from a structure theorem for -colorful graphs that exclude an arbitrary planar colorful minor. This is in stark contrast to the -colorful case provided by the Grid Theorem [76]. It turns out that, in order to fully describe the structure of -colorful graphs that exclude a fixed but arbitrary planar colorful minor, we need the full power of the so-called Graph Minor Structure Theorem (GMST) of Robertson and Seymour [72]. The definitions necessary to state the GMST are rather involved and technical, for an in-depth explanation see section 4.
We say that a graph has a -near embedding in a surface if there exists with such that such that has an embedding into with pairwise vertex-disjoint faces such that for all the ’s are pairwise vertex-disjoint for positive and for each has a path-decomposition 444A path-decomposition of a graph is a tree-decomposition where is a path. of width at most such that the vertices of appear in agreement to their cyclic ordering on the boundary of and for all The set is called the apex set, the graphs are called the vortices, and the sets are the interiors of the vortices.
@@
Notice that theorem 1.7 can also be seen as a min-max duality. This is because for big enough rainbow grids no decomposition satisfying ii) can exist. The lower bound follows from arguments of Thilikos and Wiederrecht [82].
To rephrase theorem 1.7, the absence of a rainbow grid as a colorful minor means that, with respect to any large grid-minor – or tangle (see [78, 18, 52, 38]) – in and in the absence of a big clique minor – where a local version of theorem 3.6 would hold – there is always at least one color of small bidimensionality and all vertices of such colors may be confined to the apex set and the vortices of a near-embedding. This means that for -colorful graphs, all colored vertices are fully confined to the vortices and the apex set and the entire surface part of the near embedding does not have any colored vertices. Indeed, this appears to be a far-reaching generalization of the idea to cover all colored vertices in a planar graph with few faces. We believe this provides a theoretical foundation on why many problems that are tractable on planar graphs when their annotated set is covered by few faces, remain tractable on all colorful minor-closed classes of -colorful graphs that exclude some fixed planar -colorful graph. It moreover indicates that, in algorithm design, vortices in -near embeddings may play a similar role to the one played by faces in planar graphs. It is noteworthy that theorem 1.7 is based on a much more general and technical “Local Structure Theorem” from the work of Paul, Protopapas, Thilikos, and Wiederrecht in [69] (see also [68]).
@@
It follows directly from a construction by Robertson and Seymour [76] that if a pair has the Erdős-Pósa property, then the graph must be planar. Furthermore, one may observe that if is not a colorful minor of every -segregated grid,555That is, there exists a such that is a colorful minor of every -segregated grid. then it does not have the Erdős-Pósa property. As we demonstrate in section 6, the boundary between instances that do and do not exhibit the Erdős-Pósa property shifts in subtle ways in the presence of colors – subtleties that are worth highlighting.
For every we provide the full list of graphs such that a -colorful graph has the Erdős-Pósa property if and only if it does not contain any of the graphs from as a colorful minor. In other words, we provide, for every the set of colorful minor-obstructions for the Erdős-Pósa property of -colorful graphs. We also stress that i.e., its size grows polynomially in the number of colors. Figure 3 gives a schematic definition of for every
@@
theorem 1.8 illustrates that the excluded minors for -colorful graphs do not need to have all colors. In fact contains graphs without any colored vertices and no member of has five colors no matter the choice of Our results on the Erdős-Pósa property indicate that colorful minors are a proper generalization of the concept of minors, able to express a variety of different properties.
@@
We believe that the structural theorems – theorem 1.5, theorem 1.6, and theorem 1.7 – hold significant promise for algorithm design.
Our two algorithmic results can be seen as extensions of Courcelle’s Theorem and of the main result in [81] to colorful graphs. While the structural parameters in these theorems are treewidth and Hadwiger number (that is the maximum size of a clique minor) respectively, we now introduce two parameters on colorful graphs that will serve as our new combinatorial restrictions.
The fusion of a -colorful graph is the -colorful graph where for every we have that if and only if
The strong restrictive treewidth of denoted by is defined as the restrictive treewidth of
A key condition for applying our algorithmic results is the notion of folio-representability of a class of colorful graphs (i.e., the yes-instances of some algorithmic problem), formally defined in section 8. Informally, is folio-representable if one can efficiently replace a subgraph of a colorful graph – viewed as a “protuberance”666We use the term protuberance to distinguish it from the term “protrusion” introduced in [7], where similar replacement operations were applied. Unlike protrusions, which are required to have bounded treewidth, protuberances are characterized by being “internally colorless”. with a small “boundary” interacting with the rest of the graph that has no colored vertices away from this boundary – by a bounded-size subgraph such that the resulting graph satisfies if and only if This replacement is guided by a notion of folio-equivalence: and must have the same set of rooted topological minors (up to a given size), with respect to a specified boundary Efficient computation of such replacements is enabled by results from [42], making folio-representability a natural combinatorial condition for the class
We now give the logical conditions under which our framework operates. In our first result, we require that the class be definable in Counting Monadic Second-Order Logic (CMSO); that is, there exists a formula such that if and only if models Courcelle’s Theorem says that, for every graph class definable by some CMSO-formula membership in can be decided in time We prove the following analogue for colorful graphs.

### v2

@@ Abstract
We introduce the notion of colorful minors, which generalizes the classical concept of rooted minors in graphs. A -colorful graph is defined as a pair where is a graph and assigns to each vertex a (possibly empty) subset of at most colors. The colorful minor relation enhances the classical minor relation by merging color sets at contracted edges and allowing the removal of colors from vertices. This framework naturally models algorithmic problems involving graphs with (possibly overlapping) annotated vertex sets. We develop a structural theory for colorful minors by establishing three core theorems characterizing -colorful minor-free graphs, where consists either of a clique or a grid with all vertices assigned all colors, or of grids with colors segregated and ordered on the outer face. Our results reveal that when exclusion is imposed not only on graphs but also to the way colors are distributed in them, a more refined structural landscape appears.
Leveraging our structural insights, we provide a complete classification – parameterized by the number of colors – of all colorful graphs that exhibit the Erdős–Pósa property with respect to colorful minors. On the algorithmic side, we deduce that colorful minor testing is fixed-parameter tractable. Together with the fact that the colorful minor relation forms a well-quasi-order, this implies that every colorful minor-monotone parameter on colorful graphs admits a fixed-parameter algorithm. Furthermore, we derive two algorithmic meta-theorems (AMTs) whose structural conditions are linked to extensions of treewidth and Hadwiger number on colorful graphs. Our results suggest how known AMTs can be extended to incorporate not only the structure of the input graph but also the way the colored vertices are distributed in it.
@@ 1 Introduction
Over the last decades, a prominent interplay between structural graph theory and theoretical computer science has been the search for structural properties of graphs that facilitate the design of efficient algorithms for otherwise NP-hard problems (see for example [18]). One of the most studied such parameters is a measure for the “tree-likeness” known in its most popular form as treewidth [75, 4, 5, 6]. Treewidth plays an exceptional role not only in the world of (parameterized) algorithms, but also in the theory of graph minors by Robertson and Seymour [73, 76, 77, 78]. A central theorem of Robertson and Seymour is the Grid Theorem [76] that states that any minor-closed graph class has bounded treewidth if and only if does not contain all planar graphs. This reveals a fundamental limitation in the use of treewidth (or any minor-monotone parameter really) as a parametrization for an NP-hard problem: While there exists a wide range of problems that are tractable on classes of bounded treewidth, there exist many important problems that are NP-hard even on planar graphs [34, 16]. This means that, in the regime of minor-closed graph classes, such problems are tractable if and only if a planar graph is excluded, unless .
@@
A particularly interesting class of problems tractable on graphs of bounded treewidth, but NP-hard on more general minor-closed graph classes, are certain problems whose instances come with specific vertex sets, typically called roots, terminals, or annotated vertices. A prototypical example of such a problem is Steiner Tree. The Steiner Tree problem receives a graph accompanied with a special set of annotated vertices as input and asks for the minimum number of edges in a connected subgraph of containing all vertices of The problem is known to be NP-complete on planar graphs, however it can be solved in polynomial time with respect to structural parameterizations that instead of (or in addition to) restricting the global structure of the graph (like treewidth), focus on the structure relative to the annotation [22, 25, 55], thereby providing tractable algorithms beyond the regime of minor-closed classes of bounded treewidth.
This hints at the possibility of developing a general structural and algorithmic theory for annotated graphs, providing tools that supersede the limits imposed by classical graph minor theory.
@@
Computational problems such as Steiner Tree can be naturally understood as problems on annotated graphs. Recently, several positive algorithmic results for Steiner Tree on annotated graphs excluding certain types of so-called rooted minors111An annotated graph is a rooted minor of an annotated graph if there exists a collection of pairwise vertex-disjoint connected subgraphs of such that for all there is an edge between and and for all we have have emerged [41, 46] alongside purely structural results on rooted minors [62, 44, 30] and found various applications [44, 2]. Moreover, Steiner Tree is just one example where a theory of rooted minors may open a new horizon of tractability. There exists a variety of other important problems exhibiting similar behavior, among them Multiway Cut [61, 19, 66], Disjoint paths [65, 82], #Defect Matching [17], and others [33, 16, 58]. Moreover, annotation has recently be shown to also play a key role in structural extensions of Courcelle’s Theorem [15, 13, 14] by restricting quantification to sets of low bidimensionality222The bidimensionality of a vertex set in a graph , denoted by , is the largest such that contains the -grid with all vertices annotated as a rooted minor [84]. [81].
Another area where annotation comes into play is the study of Erdős-Pósa dualities, which lies at the crossroads of structural and algorithmic graph theory. The Erdős-Pósa property for (rooted) minors asks, for a fixed (annotated) graph , whether for any (annotated) graph either contains many vertex-disjoint occurrences of as a (rooted) minor, or has a small vertex set whose deletion removes all occurrences of as a (rooted) minor. Note that not every (annotated) graph satisfies the Erdős-Pósa property. The Erdős-Pósa property has been investigated by many authors in a variety of settings [76, 68], and several key cases are also known for rooted minors [49, 70, 59, 44].
@@
While the concept of rooted minors in annotated graphs appears to be relatively well explored, a general theory for deriving results such as those above still seems to be lacking. Beyond the fact that many of the results mentioned are inspired by the theory of graph minors, no systematic extension of its powerful toolkit has yet emerged. This is somewhat surprising given the significant structural depth and algorithmic potential of these concepts, as we elaborate below.
@@ 1.1 Our results
From structure… Our core contributions are a series of structural theorems for -colorful minor-free -colorful graphs, both for general and for the particular case where is planar – under two distinct assumptions on the distribution of colors. Our results use as a departure point the notion of the bidimensionality of a vertex set and related structural concepts [68, 67, 69]. Building on our structural theorems, we derive a complete classification of all -colorful graphs that satisfy the Erdős-Pósa property, for every (under the colorful minor relation).
…to algorithms. Our next step is to apply our structural theorems to algorithmic settings, with the broader goal of developing an algorithmic graph structure theory applicable to problems involving vertex annotations. To this end, we first observe that colorful graphs are well quasi-ordered under the colorful minor relation. We also provide an algorithm that decides whether a -colorful graph is a colorful minor of a -colorful graph in time333By we mean that there is a function such that Also, given a graph we use and for the number of its vertices and edges respectively. These two facts imply that membership for any colorful minor-closed class of colorful graphs is decidable in polynomial time.
Next, we show how classical algorithmic meta-theorems (AMTs) – such as Courcelle’s Theorem and the recent algorithmic meta-theorem of Sau, Stamoulis, and Thilikos [81] – can be extended to colorful graphs, where the structural conditions are expressed via the exclusion of colorful minors. This framework enables the resolution of problems on graphs containing arbitrarily large grid or clique minors while, up to now, the exclusion of a minor appeared as a necessary (sparsity) condition.
Our results highlight a key insight:
Problem complexity does not only depend on the global structure of the input graph,
but crucially on how the annotated sets are distributed within it.
The algorithmic results presented here are not intended to be exhaustive; rather, they serve to illustrate this principle and to lay the groundwork for future exploration of algorithmic problems involving vertex annotations. In particular, we provide algorithmic applications for two of our three main structural results – Theorems˜1.2 and 1.3 – while leaving the investigation of algorithmic consequences of theorem˜1.4 to future work.
@@
Let be a non-negative integer. A -colorful graph is a pair where is a graph and444For we denote by and by For a vertex we call the palette of For illustrations, we use magenta for blue for , yellow for and green for
@@
A -colorful graph where is called restricted and a -colorful graph with is called empty. Note that any restricted -colorful graph is also a -colorful graph, so the notion of being restricted is highly dependent on the context given by the total set of colors. Finally, a -colorful graph is said to be rainbow555Notice that – in slight violation of the English language – we use the word “rainbow” as an adjective here. if for all
@@
See figure˜1 for an example. Notice that -colorful graphs are exactly annotated graphs and the notions of colorful minors and rooted minors coincide on annotated graphs.
In section˜2.4 we derive the following, as a simple consequence of a deep theorem of Robertson and Seymour [74].
@@
We prove a total of three structure theorems as follows. The first two describe the structure of graphs excluding the -colorful rainbow -clique and the -colorful rainbow -grid respectively as a colorful minor. Moreover, we generalize a theorem of Marx, Seymour, and Wollan [62], that describes the structure of -colorful graphs excluding a -grid where exactly the vertices of the first row receive color to an arbitrary number of colors.
We begin with our simplest result – excluding a rainbow clique – which can be naturally phrased in terms of a modulator to restricted graphs. For a vertex set in a graph we define the torso of in denoted by as the graph obtained from by turning the neighborhood of in into a clique for every component of In short, every -colorful graph that excludes a rainbow clique as a colorful minor has a set whose torso excludes a minor and where every component of is free of some (not necessarily the same) color.
@@
Our next theorem generalizes a theorem of Marx, Seymour, and Wollan [62] (see also [44]). Let be an integer. A -segregated grid is a -colorful graph where is the -grid where the vertices of the first column can be numbered as in order of their appearance, and we have for all and there exists a permutation of such that for every In this situation, we say that realizes . See figure˜2 for an example.
Theorem˜1.2 explains that every -colorful minor-free colorful graph can be turned into a collection of restricted graphs by deleting an -minor-free subgraph. Segregated grids capture exactly the situation where we require to be a planar graph. We may express this in form of a parameter as follows. The torso treewidth of a -colorful graph is the smallest integer such that there exists a set such that the treewidth of the torso of in is at most and for every component of the -colorful graph is restricted.
@@
contains as a colorful minor some -segregated grid and the torso treewidth of is more than , or
@@
Notice that theorem˜1.3 is optimal in the sense that the torso treewidth of segregated grids is unbounded (see lemma˜5.4). In other words, theorem˜1.3 can be seen as a mix-max duality theorem for segregated grids. We remark that for the case the torso treewidth of has been independently defined by Jansen and Swennenhuis [46] and by Hodor, La, Micek, and Rambaud [45] under different names.
It is easy to see that theorem˜1.3 is far from a structure theorem for -colorful graphs that exclude an arbitrary planar colorful minor. This is in stark contrast to the -colorful case provided by the Grid Theorem [76]. It turns out that, in order to fully describe the structure of -colorful graphs that exclude a fixed but arbitrary planar colorful minor, we need the full power of Robertson and Seymour’s Graph Minor Structure Theorem (GMST) [72, 53, 38]. The definitions necessary to state the GMST are rather involved and technical, for an in-depth explanation see section˜4.
We say that a graph has a -near embedding in a surface if there exists with such that such that has an embedding into with pairwise vertex-disjoint faces such that for all the ’s are pairwise vertex-disjoint for positive and for each has a path-decomposition 666A path-decomposition of a graph is a tree-decomposition where is a path. of width at most such that the vertices of appear in agreement to their cyclic ordering on the boundary of and for all The set is called the apex set, the graphs are called the vortices, and the sets are the interiors of the vortices.
@@
Notice that theorem˜1.4 can also be seen as a min-max duality as implied by results of Thilikos and Wiederrecht [84].
To rephrase theorem˜1.4, the absence of a rainbow grid as a colorful minor means that, with respect to any large grid-minor in and in the absence of a big clique minor – where a local version of theorem˜3.6 would hold – there is always at least one color of small bidimensionality and all vertices of such colors may be confined to the apex set and the vortices of a near-embedding. This means that for -colorful graphs, all colored vertices are fully confined to the vortices and the apex set and the entire surface part of the near embedding does not have any colored vertices. Indeed, this appears to be a far-reaching generalization of the idea to cover all colored vertices in a planar graph with few faces. We believe this provides a theoretical foundation on why many problems that are tractable on planar graphs when their annotated set is covered by few faces, remain tractable on all colorful minor-closed classes of -colorful graphs that exclude some fixed planar -colorful graph. It moreover indicates that, in algorithm design, vortices in -near embeddings may play a similar role to the one played by faces in planar graphs.
@@
It follows directly from a construction by Robertson and Seymour [76] that if a pair has the Erdős-Pósa property, then the graph must be planar. As we demonstrate in section˜6, the boundary between instances that do and do not exhibit the Erdős-Pósa property shifts in subtle ways in the presence of colors – subtleties that are worth highlighting.
For every we provide the full list of graphs such that a -colorful graph has the Erdős-Pósa property if and only if it does not contain any of the graphs from as a colorful minor. In other words, we provide, for every the set of colorful minor-obstructions for the Erdős-Pósa property of -colorful graphs. We also stress that i.e., its size grows polynomially in the number of colors. Figure˜3 gives a schematic definition of for every
@@
theorem˜1.5 illustrates that the excluded minors for -colorful graphs do not need to have all colors. In fact contains graphs without any colored vertices and no member of has five colors no matter the choice of Our results on the Erdős-Pósa property indicate that colorful minors are a proper generalization of the concept of minors, able to express a variety of different properties.
Towards a general framework for colorful parameters.
With theorem˜1.3 and theorem˜1.4 in hand, it becomes apparent that there are fundamentally distinct ways to extend a graph parameter to a colorful analogue, and that these choices can lead to remarkably different structural decompositions. Which extension should be regarded as “canonical” appears to depend strongly on the intended application and on the structural/algorithmic consequences one wishes to capture.
For instance, from the perspective of structural decompositions, torso treewidth emerges as a particularly natural analogue of treewidth, and this is further supported by the fact that it delineates the Erdős–Pósa property for colorful minors as theorem˜1.5 shows, mirroring the role played by treewidth in the uncolored setting. In contrast, when viewed through the lens of obstructions, bidimensionality provides the appropriate analogue, as it yields the correct structure theorem for colorful graphs excluding an arbitrary colorful planar graph as a colorful minor, and potentially, the correct structural parameterization to obtain precise dichotomy complexity results for problems on colorful graphs that become NP-hard in planar graphs, as previously discussed.
These observations point towards a broader, open-ended research direction, which is to better understand these phenomena by examining other classical graph parameters beyond treewidth, and by developing general principles for lifting such parameters to colorful analogues. We return to this perspective in section˜8.2 of the conclusion, where we discuss this issue in more detail and propose a conjecture.
Algorithmic results.
One of the ground-breaking results of Robertson and Seymour’s Graph Minors Series is the Graph Minor Algorithm that allows to check if a graph is a minor of a graph in time [79] (see [54, 56] for improvements). This algorithm also solves a rooted version of the Minor Checking problem which, as a special case, contains the so-called -Disjoint Paths problem.
We present a simple reduction to the rooted minor checking problem which allows us to leverage the algorithm of Korhonen, Pilipczuk, and Stamoulis [56] in order to deduce a similar result for -Colorful Minor Checking.
Theorem 1.6.
Theorems˜1.1 and 1.6 imply that the recognition problem for any proper colorful minor-closed class can be solved in polynomial time. We present below a parametric variant of this, based on the general framework introduced by Fellows and Langston [27, 28] for the existence of fixed-parameter tractable algorithms.
A (-colorful) graph parameter is a map assigning a non-negative integer to every (-colorful) graph. We say that a -colorful graph parameter is colorful minor-monotone if for every -colorful graph and every colorful minor of we have Again, combining Theorems˜1.1 and 1.6 we can deduce the following.
Theorem 1.7.
We stress that the result of theorem˜1.7 is not constructive, i.e., it does not give any way to construct the claimed algorithm, as is is based on the finiteness of the obstruction set of the class which is only existentially provided by theorem˜1.1.
@@
Our two algorithmic results can be seen as extensions of Courcelle’s Theorem and of a recent result due to Sau, Stamoulis, and Thilikos [81] to colorful graphs. While the structural parameters in these theorems are treewidth, denoted by and Hadwiger number, denoted by (that is the maximum size of a clique minor) respectively, we now introduce two parameters on colorful graphs that serve as our new combinatorial restrictions.
A key condition for applying our algorithmic results is the notion of folio-representability of a class of colorful graphs (i.e., the yes-instances of some algorithmic problem), formally defined in section˜7. Informally, is folio-representable if one can efficiently replace a subgraph of a colorful graph with colored vertices only within a small subset of the vertices of – viewed as a “protuberance”777We use the term protuberance to distinguish it from the term “protrusion” introduced in [7], where similar replacement operations were applied. Unlike protrusions, which are required to have bounded treewidth, protuberances are characterized by being “internally colorless”. interacting with the rest of the graph via the “boundary” – by a bounded-size subgraph such that the resulting graph belongs to if and only if does.
This replacement is guided by a notion of folio-equivalence: and must have the same set of rooted topological minors (up to a given size), with respect to the specified boundary Efficient computation of such replacements is enabled by results from [42], making folio-representability a natural combinatorial condition for the class
In what follows, whenever we refer to 1-colorful graphs we may alternatively use the notation instead of where . Given a graph and a set , we define the monodimensionality of in as the maximum order of a -segregated grid colorful minor of Given a -colorful graph and a color we define the monodimensionality of in as the monodimensionality of in
We now give the logical conditions under which our framework operates. In our first result, we require the class to be definable in Counting Monadic Second-Order Logic888There exists a formula such that if and only if models (CMSO). Courcelle’s Theorem says that, for every graph class definable by some CMSO-formula membership in can be decided in time We prove the following analogue for colorful graphs.